OpenStudy (anonymous):

PLEASE HELP!! What is the surface area of a sphere with radius 2? A. 4 pi units^2 B. 2 pi units^2 C. 8 pi units^2 D. 16 pi units^2

OpenStudy (anonymous):

\[4 \pi r^2\]\[4 \pi 2^2 = 4 \pi 4 = 16 \pi\]

OpenStudy (anonymous):

thanks i wanted to make sure if my answer is right thankss!

OpenStudy (anonymous):

no problem :)